# Universal Discrete-Domain Speech Enhancement

### v1

@@ IV-E4 Codec Generalization Analysis
In the UDSE framework, the neural speech codec is a crucial component, serving multiple roles including extracting global feature condition, providing clean token targets for training, and decoding the final speech output. In our experiments, we adopted DAC [45] as the neural speech codec. The coding results of clean speech from DAC serves as the upper bound for the enhancement quality achievable by UDSE. Theoretically, UDSE can also work with other RVQ-based neural speech codecs that offer comparable coding quality. To verify the generalization ability of UDSE across different neural speech codecs, we replaced the DAC with a more lightweight RVQ-based MDCTCodec [48] in UDSE (denoted by UDSE w/ MDCTCodec). The results in Table III showed no significant difference between UDSE and UDSE w/ MDCTCodec ($p=0.64$), indicating that replacing DAC with MDCTCodec did not obviously affect the performance of UDSE. This confirms that UDSE exhibits strong codec generalization to different RVQ-based codecs, laying a solid foundation for future efforts to reduce the complexity of UDSE by exploring more lightweight codec alternatives.
@@ V Conclusion
This paper proposed UDSE, a novel classification-based discrete-domain general universal SE model. Unlike conventional regression-based continuous-domain SE models that predict continuous features/waveforms, UDSE formulates the SE task as a classification problem over clean acoustic tokens quantized by a neural speech codec. The UDSE advances discrete-domain SE by introducing a self-contained framework that does not require any textual cues, semantic labels, or support from LLMs. The UDSE begins with a randomly initialized token sequence, extracts global feature conditions from degraded speech, and sequentially predicts clean tokens quantized by each VQ of a neural speech codec following the RVQ rule, where each prediction depends on the previous ones. The final clean speech is reconstructed by decoding all predicted tokens using the decoder of the neural speech codec. During training, UDSE is optimized using teacher forcing strategy with a cross-entropy classification loss. Both objective and subjective experimental results demonstrate that the proposed UDSE can effectively enhance speech affected by various single and combined distortions, confirming its strong universality and practical applicability. The introduction of UDSE not only presents a new solution to SE, but also introduces a token prediction approach that can serve as a useful reference for broader discrete-domain speech generation tasks. In future work, we aim to further optimize the UDSE architecture for improved efficiency and lightweight deployment, while expanding evaluations to cover a wider range of distortion conditions encountered in real-world applications.

### v2

@@ IV-E4 Codec Generalization Analysis
In the UDSE framework, the neural speech codec is a crucial component, serving multiple roles including extracting global feature condition, providing clean token targets for training, and decoding the final speech output. In our experiments, we adopted DAC [kumar2023high] as the neural speech codec. The coding results of clean speech from DAC serve as the upper bound for the enhancement quality achievable by UDSE. Theoretically, UDSE can also work with other RVQ-based neural speech codecs that offer comparable coding quality. To verify the generalization ability of UDSE across different neural speech codecs, we replaced the DAC with a more lightweight RVQ-based MDCTCodec [jiang2024mdctcodec] in UDSE (denoted by UDSE w/ MDCTCodec). The results in Table III showed no significant difference between UDSE and UDSE w/ MDCTCodec ($p=0.64$), indicating that replacing DAC with MDCTCodec did not obviously affect the performance of UDSE. This confirms that UDSE exhibits strong codec generalization to different RVQ-based codecs, laying a solid foundation for future efforts to reduce the complexity of UDSE by exploring more lightweight codec alternatives.
@@ V Conclusion
This paper proposed UDSE, a novel classification-based discrete-domain general universal SE model. Unlike conventional regression-based continuous-domain SE models that predict continuous features/waveforms, UDSE formulates the SE task as a classification problem over clean acoustic tokens quantized by a neural speech codec. The UDSE advances discrete-domain SE by introducing a self-contained framework that does not require any textual cues, semantic labels, or support from LLMs. The UDSE begins with a randomly initialized token sequence, extracts global feature conditions from degraded speech, and sequentially predicts clean tokens quantized by each VQ of a neural speech codec following the RVQ rule, where each prediction depends on the previous ones. The final clean speech is reconstructed by decoding all predicted tokens using the decoder of the neural speech codec. During training, UDSE is optimized using teacher forcing strategy with a cross-entropy classification loss. Both objective and subjective experimental results demonstrate that the proposed UDSE can effectively enhance speech affected by various single and combined distortions, confirming its strong universality and practical applicability. In future work, we will further optimize the model architecture and design to reduce computational complexity and latency, while continuing to explore enhancement performance under a wider range of distortion scenarios. In addition, we will explore applying the UDSE token prediction framework to other speech tasks, such as zero-shot TTS with LLMs, to extend its applicability.